# Frequency-Explicit Shape Holomorphy in Uncertainty Quantification for Acoustic Scattering

### v1

@@ 6.3 Accuracy of the PML $h$-Galerkin FEM approximation of the far-field pattern
As above, let $\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)$ be the outgoing solution of (6.6) transformed to the nominal domain. Let $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)$ be the solution of the PML variational formulation (5.1) with the corresponding right-hand side. Let $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)_{h}$ be the corresponding solution of the Galerkin equations (5.9), and let
$\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}:=C(d,k)\int\limits_{\operatorname{supp}\nabla\psi}\begin{aligned} \Big{(}\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{y}}})_{h}-\widehat{u}^{\rm inc}(\widehat{{\boldsymbol{y}}})\Big{)}\Big{(}\Delta\psi(\widehat{{\boldsymbol{y}}})-2{\rm i}k\,{\boldsymbol{q}}\cdot\nabla\psi(\widehat{{\boldsymbol{y}}})\Big{)}\cdot\\ \exp\big{(}-{\rm i}k\widehat{{\boldsymbol{y}}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}\widehat{{\boldsymbol{y}}}\end{aligned}$ (6.11)
(i.e., (6.8) with $\widehat{u}^{\rm alt}(\mathfrak{r};\widehat{{\boldsymbol{y}}})$ replaced by $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{y}}})_{h}$) denote the corresponding approximate farfield.
@@
Remark 6.5 (Comparison of Theorem 5.8 with the results of [29]).
@@
The paper [29] both
  1. 1.
    bounds the Galerkin error in terms of $\|f\|_{H^{s}}$ for arbitrary $s>0$, and
  2. 2.
A standard duality argument (see, e.g., [29, Proposition 2.1]) shows that
$\left\|{u-u_{h}}\right\|_{L^{2}}\leq C\left\|{u-u_{h}}\right\|_{H^{1}_{k}}\sup_{0\neq g\in L^{2}}\min_{v_{h}\in V_{h}}\frac{\left\|{\mathcal{S}^{*}g-v_{h}}\right\|_{H^{1}_{k}}}{\left\|{g}\right\|_{L^{2}}},$ (6.15)
and [29, Lemma 2.3] uses similar ideas to prove that
$\left|\int(u-u_{h})z\right|\leq C\left\|{u-u_{h}}\right\|_{H^{1}_{k}}\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}},$ (6.16)
If $z\in L^{2}$, then
$\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}}\leq C\Big{(}hk+(hk)^{p}k\Big{)}\left\|{z}\right\|_{L^{2}}$ (6.17)
by [51, 52], [13, Lemma 2.13], [33, Theorem 1.7]. Furthermore, if $z\in H^{p-1}$ is $k$-oscillatory, in that it satisfies the bound (5.14) (with $\widehat{f}$ replaced by $z$), then
$\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}}\leq C(hk)^{p}k\left\|{z}\right\|_{L^{2}}$ (6.18)
by [29, Part (ii) of Theorem 4.4]. The bounds (6.17) and (6.18) also hold with $\mathcal{S}^{*}$ replaced by $\mathcal{S}$ since $\mathcal{S}^{*}f=\overline{\mathcal{S}\overline{f}}$.
Therefore, if $u$ is the solution of the Helmholtz equation with $k$-oscillatory (in the sense of (5.14)) right-hand side $f\in H^{p-1}$ and $h$ is such that $(hk)^{p}k\leq C$ for sufficiently-small $C>0$, then the combination of (6.15), quasioptimality, (6.17), and (6.18) implies that
$\left\|{u-u_{h}}\right\|_{L^{2}}\leq C\Big{(}hk+(hk)^{p}k\Big{)}(hk)^{p}k\left\|{f}\right\|_{L^{2}}.$ (6.19)
Furthermore, if $z\in H^{p-1}$ is $k$-oscillatory, as in the case of the far-field pattern (6.8), then the combination of (6.16), quasioptimality, and (6.18) implies that
$\left|\int(u-u_{h})z\right|\leq C(hk)^{p}k\left\|{f}\right\|_{L^{2}}(hk)^{p}k\left\|{z}\right\|_{L^{2}}.$ (6.20)
The bound (6.19) is the same as (6.14), and the bound (6.20) is no better than (6.19) when $\|z\|_{L^{2}}\sim 1$ and when $(hk)^{p}k$ is fixed (i.e., when one chooses the least restrictive condition on $h$ allowed by the theory as $k\to\infty$). That is, in this setting of data coming from a plane-wave with the quantity of interest being the far-field pattern expressed as the linear functional (6.8), the results of [29] indeed do not give a better result than Corollary 6.4. (Note that the bounds on the error in the functional in [29, Part (ii) of Corollary 4.5 and Part (ii) of Corollary 4.6] assume that $\|z\|_{H^{s^{\prime}}}$ is independent of $k$, and thus these results have better $k$-dependence than stated above.)
@@
To this end, we assume that the shape variation is modelled as in Section 2.1, with displacement functions $r\in{\cal R}$ satisfying Assumption 2.1. Shape uncertainty will be introduced by placing a (probability) measure on a suitable set of radial displacement functions: $r=r(\omega)$, $\omega\in\Omega$, with $\Omega$ the set of elementary events in a probability space. For purposes of computation, we assume affine-parametric dependence of $r$ on a sequence ${\mathbf{y}}:=\left(y_{j}\right)_{j\in\mathbb{N}}$ of parameters. As is customary in computational UQ for PDEs (see, e.g., [59]), the measure will be constructed as product of probability measures on the co-ordinates $y_{j}$. This is to say that the dependence of $r$ on $\omega$ is expressed through a sequence $\boldsymbol{Y}=\left(Y_{j}\right)_{j\in\mathbb{N}}$ of independent identically distributed (i.i.d), $(-1,1)$-valued random variables $Y_{j}=Y_{j}(\omega)$.
@@
It is natural to assume that the shape variations of $D$ are invariant under rotations, from which we conclude $\operatorname{Cov}r({\boldsymbol{s}},{\boldsymbol{s}}^{\prime})=g({\boldsymbol{s}}\cdot{\boldsymbol{s}}^{\prime})$ for some covariance function $g:[{-1,1}]\to\mathbb{R}$. Then Karhunen-Loève expansion of $r=r(\omega)$ will yield exactly the radial spherical harmonic expansion functions $r_{j}$ of (7.2).
@@
for $p\in\mathbb{N}$.
@@ 8 Computational Shape Uncertainty Quantification
$\mathbb{E}(\breve{u}^{\rm scat}_{\infty})({\boldsymbol{q}})$ $=\int\nolimits_{{\cal P}}\breve{u}^{\rm scat}_{\infty}({{\mathbf{y}}};{\boldsymbol{q}})\,\mathrm{d}\mu({{\mathbf{y}}})\;,$ ${\boldsymbol{q}}\in\mathbb{S}^{d-1}\;,$ (8.1a)
@@
The formula (8.1a) is a so-called “ensemble average” over all admissible scatterer shapes, and involve “infinite-dimensional integrals”, whose efficient approximation by (possibly higher-order and deterministic) quadrature formulas is addressed next.
@@ 8.1 High-Dimensional Smolyak / Sparse-Grid Quadrature
The QoI (8.1a) being a countably-parametric, deterministic integral, we discuss two classes of numerical integration: first, so-called Smolyak-Quadrature (see [66] and references there) and second, so-called Higher-Order QMC Quadrature, in particular so-called interlaced polynomial lattice rules (“IPL-QMC integration”, see [26]). In these references, convergence rates of both numerical integration methods have been proved to be independent of the dimension of the domain of integration, under suitable quantified holomorphy of the integrands. In view of the preceding discussion on quantified, wavenumber-explicit holomorphy of the parametric solutions $\{\breve{{\mathbf{y}}}\mapsto\hat{u}(\breve{{\mathbf{y}}}):\breve{{\mathbf{y}}}\in{\cal P}\}$, we connect the parametric holomorphy analysis to [26, 66]. We start by recalling a suitable concept of quantified parametric holomorphy, which implies dimension-independent convergence rates of (i) Smolyak-quadratures [65, 66], (ii) sparse-grid interpolants [14, 65] and (iii) higher-order Quasi-Monte Carlo quadratures [26].
8.1.1 $({\boldsymbol{b}},\bar{p},\varepsilon)$-Holomorphy
We quantify the domain of parametric holomorphy of the (analytic continuation of the) affine-parametric function $\{\breve{{\mathbf{y}}}\in{\cal P}\mapsto u_{\breve{{\mathbf{y}}}}:\breve{{\mathbf{y}}}\in{\cal P}\}$. We still work with the $k^{-1}$-scaling and the affine-parametric parametrizations $\breve{{\mathbf{y}}}\mapsto\mathfrak{r}(\breve{{\mathbf{y}}})$ from (7.4) and with the globally supported radial expansion functions $r_{j}$ from (7.2).
We present quantified parametric holomorphy for a generic “target” Banach space $X$. Observe that the QoI’s in (8.1a) take values in the separable, complex Banach space $X\in\{C^{0}(\mathbb{S}^{d-1};\mathbb{C}),L^{2}(\mathbb{S}^{d-1};\mathbb{C})\}$.
@@ 8.1.1 $({\boldsymbol{b}},\bar{p},\varepsilon)$-Holomorphy
The significance of $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy lies in the fact that holomorphic maps between (complex) Banach spaces become, upon adopting affine-parametric representation of their arguments w.r. to a suitable representation system (such as, e.g. $\{\psi_{j}\}_{j\geq 1}$ with $\psi_{j}\sim\beta_{j}r_{j}$ where $r_{j}$ is as in Remark 7.2) $({\boldsymbol{b}},\varepsilon)$-holomorphic maps in terms of the coefficient sequences in the representation of inputs (see [66, Lem. 3.3]). The summability exponent $\bar{p}\in(0,1)$ of the sequence ${\boldsymbol{b}}$ determines the convergence rate of suitable Smolyak quadratures in (8.8) below.
8.1.2 Estimating the Smolyak Quadrature Error
We recall the definition of a sparse-grid Smolyak quadrature: given a sequence $(\chi_{n;j})_{j=0}^{n}$, $n\in\mathbb{N}_{0}$ of $n$-tuples of pairwise distinct points in $[-1,1]$, Smolyak quadratures are built on corresponding univariate interpolatory quadrature rules $(Q_{n})_{n\geq 0}$ with nodes $\chi_{n;0},...,\chi_{n;n}\subset[-1,1]$ and corresponding weights $w_{n;j}>0$ w.r. to the uniform (probability) measure $\frac{1}{2}\lambda^{1}$, i.e.
@@ 8.1.2 Estimating the Smolyak Quadrature Error
By construction, $Q_{n}$ is exact for univariate polynomials of degree $n$. The weights $w_{n;j}$ can be negative, in general. For the error bound [66, Theorem 2.16] to hold, we assume stability of the univariate quadrature points in the sense that there exist constants $C,\vartheta>0$ such that
$\forall n\in\mathbb{N}_{0}:\quad\sup_{0\neq f\in C^{0}([-1,1])}\frac{|Q_{n}f|}{\|f\|_{C^{0}([-1,1])}}\leq C(n+1)^{\vartheta}$ (8.4)
@@
Multivariate anisotropic quadratures are built from the univariate hierarchy $(Q_{n})_{n\geq 0}$ by tensorization. Let =${\mathcal{F}}=\{{\boldsymbol{\nu}}\in\mathbb{N}_{0}^{\mathbb{N}}:|{\boldsymbol{\nu}}|<\infty\}$ denote the set of finitely supported multiindices, and let $\Lambda\subset{\mathcal{F}}$ be a downward closed777 We recall (e.g. [15, Def. 1.1]) that an index set $\Lambda\subset{\mathcal{F}}$ is downward closed (“d.c.” for short) if ${\boldsymbol{\nu}}\in\Lambda$ and ${\boldsymbol{\mu}}\leq{\boldsymbol{\nu}}$ implies ${\boldsymbol{\mu}}\in\Lambda$. Here, ${\boldsymbol{\mu}}\leq{\boldsymbol{\nu}}$ means $\mu_{j}\leq\nu_{j}$ for all $j$. finite index set. Then for ${\boldsymbol{\nu}}\in{\mathcal{F}}$, we define the multivariate tensor product quadrature $Q_{\boldsymbol{\nu}}:=\bigotimes_{j\in\mathbb{N}}Q_{\nu_{j}}$. Then the Smolyak Quadrature for a d.c. set $\Lambda\subset{\mathcal{F}}$ is defined by
@@
We refer to [66, Section 2.2] for details.
The main result from [66, Section 2.5] on the convergence rate of $Q_{\Lambda}$ for suitable downward closed sets $\Lambda\subset{\mathcal{F}}$ of “active quadrature orders” ${\boldsymbol{\nu}}\in\Lambda$ is as follows.
@@
    [$k$-independent parametric holomorphy] The parametric maps $U$ in (8.7) are $({\boldsymbol{b}},p,\varepsilon)$ holomorphic in the sense of Definition 8.1, with some $\varepsilon>0$ independent of $k$, and
@@
This is [66, Thm. 2.16 and Thm.  4.3, Eqn.(4.1)], with the $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy followwing from [66, Lemma 3.3]. Inspection of the proofs in [66] reveals, in particular, that the constant $C>0$ in the quadrature error bound in the statement of [66, Thm.  4.3, Eqn.(4.1)] scales linearly in the integrand modulus $M_{u}$, whence the bound (8.8).
Here, as in (8.1a), the measure $\mu$ in (8.8) is a countable product probability measure obtained from the univariate scaled Lebesgue measure, $\frac{1}{2}\lambda^{1}$. For nested collections of univariate integration points in the sense of [66, Def. 2.1] it holds that $|{\mathrm{pts}}(\Lambda,{\boldsymbol{\chi}})|=|\Lambda|$ ([66, Lemma 2.1]),

### v2

@@ 6.3 Accuracy of the PML $h$-Galerkin FEM approximation of the far-field pattern
As above, let $\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)$ be the outgoing solution of (6.6) transformed to the nominal domain. Let $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)$ be the solution of the PML variational formulation (5.1) with the corresponding right-hand side. Let $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)_{h}$ be the corresponding solution of the Galerkin equations (5.9), and let, for ${\boldsymbol{q}}\in\mathbb{S}^{d-1}$,
$\boxed{\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}:=C(d,k)\int\limits_{\operatorname{supp}\nabla\psi}\begin{aligned} \Big{(}\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{y}}})_{h}-\widehat{u}^{\rm inc}(\widehat{{\boldsymbol{y}}})\Big{)}\Big{(}\Delta\psi(\widehat{{\boldsymbol{y}}})-2{\rm i}k\,{\boldsymbol{q}}\cdot\nabla\psi(\widehat{{\boldsymbol{y}}})\Big{)}\cdot\\ \exp\big{(}-{\rm i}k\widehat{{\boldsymbol{y}}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}\widehat{{\boldsymbol{y}}}\end{aligned}}$ (6.11)
(i.e., (6.8) with $\widehat{u}^{\rm alt}(\mathfrak{r};\widehat{{\boldsymbol{y}}})$ replaced by $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{y}}})_{h}$) denote the corresponding approximate far-field.
@@
Remark 6.5.
@@
In §A we describe how, given the current state-of-the-art FEM convergence theory, we cannot do better than arguing as in Points 1 and 2 above.
@@
To this end, we further constrain the generic shapes that were introduced in Section 2.1, with displacement functions $r\in{\cal R}$ satisfying Assumption 2.1. We adopt a affine-parametric representation of the scatterers’ shape. Probabilistic models of shape uncertainty will subseuently be introduced by placing a (probability) measure on the affine parameter sequences which occur in the radial displacement functions, yielding $r=r(\omega)$, $\omega\in\Omega$, with $\Omega$ the set of elementary events in a probability space. Specifically, we assume affine-parametric dependence of $r$ on a sequence ${\mathbf{y}}:=\left(y_{j}\right)_{j\in\mathbb{N}}$ of parameters. As is customary in computational UQ for PDEs (see, e.g., [62, 48]), the measure will be constructed as product of probability measures on the co-ordinates $y_{j}$. This is to say that the dependence of $r$ on $\omega$ is expressed through a sequence $\boldsymbol{Y}=\left(Y_{j}\right)_{j\in\mathbb{N}}$ of independent identically distributed (i.i.d), $(-1,1)$-valued random variables $Y_{j}=Y_{j}(\omega)$.
@@
In applications where $D$ is, for example, an “imperfect sphere” [61, Chapter 1] one may assume that the shape variations of $D$ are invariant under rotations, from which we conclude $\operatorname{Cov}r({\boldsymbol{s}},{\boldsymbol{s}}^{\prime})=g({\boldsymbol{s}}\cdot{\boldsymbol{s}}^{\prime})$ for some covariance function $g:[{-1,1}]\to\mathbb{R}$. Then Karhunen-Loève expansion of $r=r(\omega)$ will yield exactly the radial spherical harmonic expansion functions $r_{j}$ of (7.2).
@@
for $p\in\mathbb{N}$ (as introduced in the paragraph following (5.9)). Examples of admissible sequences $\left(\beta_{j}\right)_{j}$ for particular choices of the geometry variations $r_{j}$ will be presented in Section 8.2, see (8.11).
@@ 8 Computational Shape Uncertainty Quantification
$\begin{array}[]{rl}\mathbb{E}(\breve{u}^{\rm scat}_{\infty})({\boldsymbol{q}})&=\int\nolimits_{{\cal P}}\breve{u}^{\rm scat}_{\infty}({{\mathbf{y}}};{\boldsymbol{q}})\,\mathrm{d}\mu({{\mathbf{y}}})\;,\quad{\boldsymbol{q}}\in\mathbb{S}^{d-1}\;,\end{array}$ (8.1)
@@
The formula (8.1) is a so-called “ensemble average” over all admissible scatterer shapes, and involve “infinite-dimensional integrals”, whose efficient approximation by (possibly higher-order and deterministic) quadrature formulas is addressed next.
@@ 8.1 High-Dimensional Smolyak / Sparse-Grid Quadrature
The QoI (8.1) being a countably-parametric, deterministic integral, we discuss two classes of numerical integration: first, so-called Smolyak-Quadrature (see [69] and references there) and second, so-called Higher-Order QMC Quadrature, in particular so-called interlaced polynomial lattice rules (“IPL-QMC integration”, see [28]). In these references, convergence rates of both numerical integration methods have been proved to be independent of the dimension of the domain of integration, under suitable quantified holomorphy of the integrands. In view of the preceding discussion on quantified, wavenumber-explicit holomorphy of the parametric solutions $\{\breve{{\mathbf{y}}}\mapsto\hat{u}(\breve{{\mathbf{y}}}):\breve{{\mathbf{y}}}\in{\cal P}\}$, we connect the parametric holomorphy analysis to [28, 69]. We start by recalling a suitable concept of quantified parametric holomorphy, which implies dimension-independent convergence rates of (i) Smolyak-quadratures [68, 69], (ii) sparse-grid interpolants [16, 68] and (iii) higher-order Quasi-Monte Carlo quadratures [28].
In order to quantify the holomorphic parameter dependence, we quantify the domain of parametric holomorphy of the (analytic continuation of the) affine-parametric function $\{\breve{{\mathbf{y}}}\in{\cal P}\mapsto u_{\breve{{\mathbf{y}}}}:\breve{{\mathbf{y}}}\in{\cal P}\}$. We still work with the $k^{-1}$-scaling and the affine-parametric parametrizations $\breve{{\mathbf{y}}}\mapsto\mathfrak{r}(\breve{{\mathbf{y}}})$ from (7.4) and with the globally supported radial expansion functions $r_{j}$ from (7.2).
We present quantified parametric holomorphy for a generic “target” Banach space $X$. Observe that the QoI’s in (8.1) take values in the separable, complex Banach space $X\in\{C^{0}(\mathbb{S}^{d-1};\mathbb{C}),L^{2}(\mathbb{S}^{d-1};\mathbb{C})\}$.
@@
The significance of $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy lies in the fact that holomorphic maps between (complex) Banach spaces become, upon adopting affine-parametric representation of their arguments w.r. to a suitable representation system (such as, e.g. $\{\psi_{j}\}_{j\geq 1}$ with $\psi_{j}\sim\beta_{j}r_{j}$ where $r_{j}$ is as in Remark 7.2) $({\boldsymbol{b}},\varepsilon)$-holomorphic maps in terms of the coefficient sequences in the representation of inputs (see [69, Lemma 3.3]). The summability exponent $\bar{p}\in(0,1)$ of the sequence ${\boldsymbol{b}}$ determines the convergence rate of suitable Smolyak quadratures in (8.8) below.
Next, we proceed to estimating the Smolyak Quadrature Error. To this end, we recall the definition of a sparse-grid Smolyak quadrature: given a sequence $(\chi_{n;j})_{j=0}^{n}$, $n\in\mathbb{N}_{0}$ of $n$-tuples of pairwise distinct points in $[-1,1]$, Smolyak quadratures are built on corresponding univariate interpolatory quadrature rules $(Q_{n})_{n\geq 0}$ with nodes $\chi_{n;0},...,\chi_{n;n}\subset[-1,1]$ and corresponding weights $w_{n;j}>0$ w.r. to the uniform (probability) measure $\frac{1}{2}\lambda^{1}$, i.e.
@@
By construction, $Q_{n}$ is exact for univariate polynomials of degree $n$. The weights $w_{n;j}$ can be negative, in general. For the error bound [69, Theorem 2.16] to hold, we assume stability of the univariate quadrature points [69, Eqn. (2.3)], i.e. in the sense that there exist a constant $\vartheta>0$ such that
$\forall n\in\mathbb{N}_{0}:\quad\sup_{0\neq f\in C^{0}([-1,1])}\frac{|Q_{n}f|}{\|f\|_{C^{0}([-1,1])}}\leq(n+1)^{\vartheta}$ (8.4)
@@
Multivariate anisotropic quadratures are built from the univariate hierarchy $(Q_{n})_{n\geq 0}$ by tensorization. Let =${\mathcal{F}}=\{{\boldsymbol{\nu}}\in\mathbb{N}_{0}^{\mathbb{N}}:|{\boldsymbol{\nu}}|<\infty\}$ denote the set of finitely supported multiindices, and let $\Lambda\subset{\mathcal{F}}$ be a downward closed888 We recall (e.g. [16, Definition 1.1]) that an index set $\Lambda\subset{\mathcal{F}}$ is downward closed (“d.c.” for short) if ${\boldsymbol{\nu}}\in\Lambda$ and ${\boldsymbol{\mu}}\leq{\boldsymbol{\nu}}$ implies ${\boldsymbol{\mu}}\in\Lambda$. Here, ${\boldsymbol{\mu}}\leq{\boldsymbol{\nu}}$ means $\mu_{j}\leq\nu_{j}$ for all $j$. finite index set. Then for ${\boldsymbol{\nu}}\in{\mathcal{F}}$, we define the multivariate tensor product quadrature $Q_{\boldsymbol{\nu}}:=\bigotimes_{j\in\mathbb{N}}Q_{\nu_{j}}$. Then the Smolyak Quadrature for a d.c. set $\Lambda\subset{\mathcal{F}}$ is defined by
@@
We refer to [69, Section 2.2] for details.
The main result from [69, Section 2.5] on the convergence rate of $Q_{\Lambda}$ for suitable downward closed sets $\Lambda\subset{\mathcal{F}}$ of “active quadrature orders” ${\boldsymbol{\nu}}\in\Lambda$ is as follows.
@@
    [$k$-independent parametric holomorphy] The parametric maps $U$ in (8.7) are $({\boldsymbol{b}},\bar{p},\varepsilon)$ holomorphic in the sense of Definition 8.1, with some $\varepsilon>0$ independent of $k$, and
@@
This is [69, Theorem 2.16 and Theorem  4.3, Eqn.(4.1)], with the $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy following from [69, Lemma 3.3]. Inspection of the proofs in [69] reveals, in particular, that the constant $C>0$ in the quadrature error bound in the statement of [69, Theorem  4.3, Eqn.(4.1)] scales linearly in the integrand modulus $M_{u}$, whence the bound (8.8).
Here, as in (8.1), the measure $\mu$ in (8.8) is a countable product probability measure obtained from the univariate scaled Lebesgue measure, $\frac{1}{2}\lambda^{1}$. For nested collections of univariate integration points in the sense of [69, Definition 2.1] it holds that $|{\mathrm{pts}}(\Lambda,{\boldsymbol{\chi}})|=|\Lambda|$ ([69, Lemma 2.1]),